Let G be the set of positive real numbers except 1. Define α∗β = α. Then:

(a) show that (G,∗) is a group.

(b) Is G abelian? if not, find its center.

(c) Give an automorphism of G. 


1
Expert's answer
2021-08-06T07:48:39-0400

"\\text{Let $\\alpha,\\beta,\\gamma \\in G$}\\\\ \\text{Consider }\\\\\n\n\n\u03b1\u2217\u03b2=\u03b1^{ln\u2061\u03b2} \\\\\n\n\\text{Clearly, \u03b1\u2217\u03b2 is positive since the power of any positive real number is always positive. } \\\\\n\n\\text{Next, suppose \u03b1\u2217\u03b2=1. Then we have}\\\\\n\\alpha^{\\ln \\beta}=1=\\alpha^0 \\Rightarrow \\ln \\beta =0 \\\\ \\Rightarrow \\beta=1\\\\\n\n\\text{which is a contradiction. Hence $\\alpha * \\beta \\in G$} \\\\\n\\text{Next, consider }\\\\ \\alpha * (\\beta * \\gamma)=\\alpha^{\\ln\\left(\\beta^{\\ln\\gamma}\\right)}=\\alpha^{(\\ln\\gamma)(\\ln\\beta)} \\\\ =\\alpha^{(\\ln\\beta)(\\ln\\gamma)}=\\left(\\alpha^{(\\ln\\beta)}\\right)^{\\ln\\gamma} \\\\=(\\alpha * \\beta) * \\gamma\\\\\n\n\\text{Next, consider}\\\\\n\\alpha * e=\\alpha^{\\ln e}=\\alpha^1=\\alpha\\\\\n\n\\text{Similarly, $e*\\alpha=e^{\\ln\\alpha}=\\alpha$.} \\\\\n\n\\text{Finally, consider }\\\\\n\\alpha*e^{\\frac{1}{\\ln\\alpha}}=\\alpha^{\\ln{(e^{\\frac{1}{\\ln\\alpha}})}}=\\alpha^{\\frac{1}{\\ln\\alpha}} \\\\=\\alpha^{log_{\\alpha}e}=e \\\\\n\n\\text{Similarly, }\\\\\ne^{\\frac{1}{\\ln\\alpha}}*\\alpha=\\left(e^{\\frac{1}{\\ln\\alpha}}\\right)^{\\ln\\alpha}=e^{\\frac{\\ln\\alpha}{\\ln\\alpha}} \\\\=e^1=e \\\\\n\\text{So the inverse of $\\alpha \\in G$ is $e^{\\frac{1}{\\ln\\alpha}}$}\u200b.\n\\\\\n\\text{Since G satisfies all the conditions of a group, we therefore conclude that G is a group.} \\\\\n\n\n\\text{Consider $e^{(\\ln\\alpha)(\\ln\\beta)}$. We can write it as} \\\\\ne^{(\\ln\\alpha)(\\ln\\beta)}=\\left(e^{(\\ln\\alpha)}\\right)^{(\\ln\\beta)}=\\alpha^{\\ln\\beta} \\\\=\\alpha*\\beta\\\\\n\n\\text{and} \\\\\ne^{(\\ln\\alpha)(\\ln\\beta)}=e^{(\\ln\\beta)(\\ln\\alpha)} \\\\=\\left(e^{(\\ln\\beta)}\\right)^{(\\ln\\alpha)}=\\beta^{\\ln\\alpha}=\\beta*\\alpha \\\\\n\n\\text{So $G$ is abelian.}\\\\\n\n\n\\text{Define a function $i:G\\rightarrow G$ such that $i(g)=g\\, \\forall g\\in G$}\\\\\n\n\n\\text{Let $g,h\\in G$. Suppose $i(g)=i(h)$. Then $g=h$. So iii is one-to-one.} \\\\\n\n\n\\text{Suppose $g\\in G$ . Then $g=i(g)$. So iii is onto.}\n\n\n\\text{Finally, consider}\\\\\ni(g)*i(h)=g*h=i(g*h)\\\\\n\\text{Hence iii is an isomorphism.}\n\n\\\\\n\\text{Thus we have an automorphism of G.}"